# MoLoRec: A Generalizable and Efficient Framework for LLM-Based Recommendation

### v1

@@ A.2. Baselines
We compare MoLoRec with traditional recommendation methods (BPR-MF, GRU4Rec, SASRec, and FMLP-Rec), transferable sequential recommenders (UniSRec, VQ-Rec), LLM-based recommenders (Qwen2-7B, RecFormer, P5, TALLRec) and our ablation counterparts (MoLoRec-G, MoLoRec-S).
BPR-MF (Rendle et al., 2009) is one of the most representative collaborative filtering models.
@@
GRU4Rec (Hidasi et al., 2016) is a seminal method that uses RNNs to model user action sequences for session-based recommendation.
@@
SASRec (Kang and McAuley, 2018) is a representative sequential recommender model that adopts a self-attention mechanism to learn the item dependency from user interactions.
@@
FMLP-Rec (Zhou et al., 2022) is an all-MLP model with learnable filters for sequential recommendation tasks.
@@
UniSRec (Hou et al., 2022) equips textual item representations with an MoE-enhanced adaptor for domain fusion and adaptation. Both item-sequence and sequence-sequence contrastive learning tasks are designed for pre-training transferable sequence representations.
@@
VQ-Rec (Hou et al., 2023) learns vector-quantized item representations for transferable sequential Recommenders.
@@
Qwen2-7B333https://huggingface.co/Qwen/Qwen2-7B-Instruct is a well-known open-source LLM. In our experiments, we choose it as MoLoRec ’s LLM backbone.
@@
P5 (Geng et al., 2022) is a unified LLM-based recommendation framework. It is built on T5 by fine-tuning with multiple recommendation tasks.
@@
TALLRec (Bao et al., 2023) learns the recommendation task based on prompts consisting solely of text and fine-tunes the LLMs using the LoRA.
@@
MoLoRec-G is an ablation counterpart of our proposed framework. It only underwent stage 1, utilizing only the domain-general LoRA module.
@@
MoLoRec-S is an ablation counterpart of our proposed framework. It only underwent stage 2, utilizing only the domain-general LoRA module.
A.3. Implementation Details
@@
To ensure a fair comparison, the experimental settings are standardized as follows: For traditional recommendation methods (BPR-MF, GRU4Rec, SASRec, and FMLP-Rec), the learning rate is set to 0.001, and the Adam optimizer is employed. The batch size is set to 256, and the embedding dimension is set to 64. Regarding transferable sequential recommenders (UniSRec, VQ-Rec), these models utilize a BERT for text processing. Specifically, the pre-trained models provided by the original authors are fine-tuned on our dataset. For RecFormer, the pre-trained model provided by the original work is also fine-tuned on the downstream tasks. In the case of P5 and TALLRec, an identical instruction fine-tuning template is used to align the original models with the recommendation tasks. Here, P5 undergoes full fine-tuning, while TALLRec is fine-tuned using a LoRA approach with a rank of 16. This setup standardizes the evaluation framework across different recommendation methodologies, ensuring comparability and fairness in assessing their performance. We use Qwen2-7B as the LLM backbone for MoLoRec. For parameter-efficient finetuning(PEFT) conducted on NVIDIA RTX 4090(24G) GPUs, we adopt low-rank adaption(LoRA) with LoRA rank as 16, LoRA alpha as 32, and LoRA dropout as 0.05 to get general LoRA adapter and target LoRA adapter. The learning rate is selected from 1e-4,2e-4 and the batch size is set to 128. For LoRA adapters fusion weights learning conducted on two NVIDIA RTX 4090(24G) GPUs, the batch size is set to 60, number of test samples for training is selected from 50,100. The number of tokens at the beginning of each title involved in training is set to 3 for Toys and Sports and 5 for Beauty because titles in Beauty are longer than Toys and Sports. In order to reduce GPU memory usage, we employed gradient checkpointing techniques. Since the model structure has not changed, we use the VLLM inference acceleration framework to perform inference and then evaluate the results. Other implementation details are available in our open-source code.

### v2

@@ A.2 Baselines
We compare RecCocktail with traditional recommendation methods (BPR-MF, GRU4Rec, SASRec, and FMLP-Rec), transferable sequential recommenders (UniSRec, VQ-Rec), LLM-based recommenders (Qwen2-7B-zeroshot, RecFormer, P5, TALLRec, AlphaRec) and our ablation counterparts (RecCocktail-G, RecCocktail-S).
BPR-MF (Rendle et al. 2009) is one of the most representative collaborative filtering models.
@@
GRU4Rec (Hidasi et al. 2016) is a seminal method that uses RNNs to model user action sequences for session-based recommendation.
@@
SASRec (Kang and McAuley 2018) is a representative sequential recommender model that adopts a self-attention mechanism to learn the item dependency from user interactions.
@@
FMLP-Rec (Zhou et al. 2022) is an all-MLP model with learnable filters for sequential recommendation tasks.
@@
UniSRec (Hou et al. 2022) equips textual item representations with an MoE-enhanced adaptor for domain fusion and adaptation. Both item-sequence and sequence-sequence contrastive learning tasks are designed for pre-training transferable sequence representations.
@@
VQ-Rec (Hou et al. 2023) learns vector-quantized item representations for transferable sequential Recommenders.
@@
Qwen2-7B-zeroshot333https://huggingface.co/Qwen/Qwen2-7B-Instruct is a well-known open-source LLM. In our experiments, we choose it as RecCocktail ’s LLM backbone.
RecFormer (Li et al. 2023) models user preferences and item features using the LongFormer (Beltagy, Peters, and Cohan 2020) backbone, transforming sequential recommendation into a task of predicting the next item as if predicting the next sentence, by converting item attributes into a sentence format.
@@
P5 (Geng et al. 2022) is a unified LLM-based recommendation framework. It is built on T5 by fine-tuning with multiple recommendation tasks.
@@
TALLRec (Bao et al. 2023) learns the recommendation task based on prompts consisting solely of text and fine-tunes the LLMs using the LoRA. We use LLaMA-7B as the LLM backbone, following the original paper. For constructing instruction data, we adopt the same prompt format as us.
AlphaRec (Sheng et al. 2025) integrates language representations for recommendation. It can adjust recommendations according to text-based user intentions, enabling recommenders to evolve into intention-aware systems through a straightforward paradigm shift.
RecCocktail-WA is an ablation counterpart of our proposed framework. It uses Weight Averaging (WA) to merge domain-general and domain-specific LoRAs instead of the Entropy-Guided Adaptive Merging strategy.
@@
RecCocktail-G is an ablation counterpart of our proposed framework. It only underwent stage 1, utilizing only the domain-general LoRA module.
@@
RecCocktail-S is an ablation counterpart of our proposed framework. It only underwent stage 2, utilizing only the domain-specific LoRA module.
A.3 Implementation Details
@@
To ensure a fair comparison, the experimental settings are standardized as follows: For traditional recommendation methods (BPR-MF, GRU4Rec, SASRec, and FMLP-Rec), the learning rate is set to 0.001, and the Adam optimizer is employed. The batch size is set to 256, and the embedding dimension is set to 64. Regarding transferable sequential recommenders (UniSRec, VQ-Rec), these models utilize a BERT for text processing. Specifically, the pre-trained models provided by the original authors are fine-tuned on our dataset. For RecFormer, the pre-trained model provided by the original work is also fine-tuned on the downstream tasks. In the case of P5 and TALLRec, an identical instruction fine-tuning template is used to align the original models with the recommendation tasks. Here, P5 undergoes full fine-tuning, while TALLRec is fine-tuned using a LoRA approach with a rank of 16. This setup standardizes the evaluation framework across different recommendation methodologies, ensuring comparability and fairness in assessing their performance. We use Qwen2-7B as the LLM backbone for RecCocktail. For PEFT conducted on NVIDIA RTX 4090(24G) GPUs, we adopt LoRA with rank as 16, alpha as 32, and LoRA dropout as 0.05 to get general LoRA and specific LoRA. The learning rate is selected from {1e-4, 2e-4} and the batch size is set to 128. For LoRA adapters fusion weights learning conducted on two NVIDIA RTX 4090(24G) GPUs, the batch size is set to 60, the number of test samples for learning and is grid searched from {50, 100}. The number of tokens at the beginning of each title involved in training is set to 3 for Toys and Sports and 5 for Beauty because titles in Beauty are longer than Toys and Sports. In order to reduce GPU memory usage, we employed gradient checkpointing techniques. Since the model structure has not changed, we use the VLLM inference acceleration framework to perform inference and then evaluate the results. Other implementation details are available in our source code.
A.4 Experiments on More LLM Backbones
To evaluate the robustness of the RecCocktail framework to different LLM backbones, we also utilize Llama-3.1-8B444https://huggingface.co/meta-llama/Llama-3.1-8B as LLM backbone and observe the experimental results. As shown in Table 5, when switching the LLM backbone to LLaMA-3.1-8B, RecCocktail improves over both RecCocktail-G and RecCocktail-S on all datasets, indicating the effectiveness of adaptive merging.
[…]
A.5 Case Study
We further conduct a case study to delve deeper into the recommendation results of RecCocktail. We randomly selected a user from the Movielens-1M test set, provided their historical viewing records and a candidate set, and asked both our model and ChatGPT to make movie recommendations along with explanations for their choices. The outputs are shown in Figure 5. We find that RecCocktail successfully generalizes to the explainable recommendation task. It accurately captured the user’s preference for action movies from their historical viewing records and leveraged world knowledge to provide an accurate interpretation of the plot of Die Hard. In contrast, GPT-4 lacks domain-specific knowledge in the recommendation, incorrectly associating the action movie Die Hard with science fiction films like Star Wars, Alien, and Terminator, resulting in unreasonable explanation outcomes. This case study further highlights RecCocktail’s task generalization capability and its deep understanding of recommendation knowledge.
Appendix B Related Work
Sequential Recommendation. Sequential recommendation utilizes users’ interaction histories to predict the next relevant item. Deep learning-based methods, such as RNNs (Hidasi et al. 2016; Li et al. 2017), GNNs (Qiu et al. 2019; Chang et al. 2021), and attention mechanisms (Kang and McAuley 2018; Zhou et al. 2020; Xie et al. 2022), have become mainstream but rely solely on item IDs, limiting their adaptability to new scenarios. To improve transferability, transferable sequential recommendation studies (Hou et al. 2022, 2023; Li et al. 2023; Yuan et al. 2023) explore leveraging textual features to enhance item representations. These approaches improve the transferability and robustness of recommender systems. Nowadays, LLMs offer new opportunities for sequential modeling, promising more robust and generalizable recommender systems.
LLM-Based Recommendation. With the rise of LLMs, interest in LLM-based recommender systems has grown, leveraging LLMs as core engines. Early studies (Dai et al. 2023; Sanner et al. 2023; Wang et al. 2023) explore their zero-shot/few-shot potential via in-context learning (Dong et al. 2024). However, the gap between LLMs’ pretraining on general text and recommendation-specific needs leads to suboptimal performance. To address this, recent research follows two paradigms: breadth-oriented and depth-oriented. The former integrates multi-domain (Tang et al. 2024; Peng et al. 2024) or multi-task (Geng et al. 2022; Zhang et al. 2024; Cui et al. 2022; Peng et al. 2024) recommendation data to construct extensive recommendation world knowledge, paving the way for developing a generalizable LLM-based recommender. For example, Peng et al.(Peng et al. 2024) build large-scale e-commerce instruction dataset ECInstruct and develop generalist LLM for e-commerce. P5 (Geng et al. 2022) designs prompts to unify 5 recommendation tasks and presents a unified text-to-text recommendation paradigm. Depth-oriented paradigm seeks to enable LLMs to deeply comprehend recommendation tasks within specific domains. Key areas of focus include: 1) the in-depth alignment of domain-specific recommendation knowledge, such as collaborative signals (Lin et al. 2024a; Kim et al. 2024a; Liao et al. 2024; Ren et al. 2024; Kong et al. 2024). The introduction of collaborative signals effectively improves model performance in warm-start scenarios. However, this comes at the cost of reduced generalization, making it challenging to adapt to new domains and cold-start situations. Another research focus is 2) the development of efficient alignment methods between large language models and recommendation tasks. These methods include leveraging (Bao et al. 2023; Kong et al. 2024) LoRA fine-tuning techniques and designing data-efficient fine-tuning strategies (Lin et al. 2024b; Zheng et al. 2024). In fact, the two paradigms have complementary advantages. In this work, we investigate how to integrate the advantages of both paradigms to simultaneously achieve both breadth and depth.
Model Merging. Model merging aims to combine multiple expert models into a more capable single model, offering benefits such as reduced storage and serving costs (Yang et al. 2024b). Model merging techniques are applied in various scenarios such as unlearning old-knowledges in LLMs (Zaman, Choshen, and Srivastava 2024), understanding content across multiple modalities (Aiello et al. 2024; Chen et al. 2024), generating images with different styles or achieving image-style transformation (Biggs et al. 2024). Previous attempts involve merging multiple models, all initially trained on the same task, with the aim of enhancing the model’s overall generalization (Gupta, Serrano, and DeCoste 2020; Wang et al. 2022). Inspired by these works, we apply model merging to the LoRA modules, leveraging their ability to integrate domain-general and domain-specific knowledge effectively.